# Stochastic Volatility in Mean: Efficient Analysis by a Generalized Mixture Sampler

### v1

@@ 6.2 Estimation results for SVM model
Using the same prior distributions as in illustrative examples in Section 4, the proposed SVM models with leverage are fitted to TB, DGS and AAA data. We iterated MCMC simulation 150,000 times after discarding initial 30,000 MCMC draws as burn-in period using Algorithm 3 in Appendix B. The acceptance rates of the MH algorithms for $\alpha$ and $(\alpha,h)$ are 70.2% and 20.7% with TB data, 69.5% and 13.2% with DGS data, and 51.1% and 36.6% with AAA data, respectively.
@@
$\mu$1.882 0.623 (-3.163, -0.656) 109 0.003
-3.824 0.641 (-5.053, -2.319) 130 0.000
3.699 0.882 ( 2.136, 5.877) 86 1.000
$\phi$ 0.920 0.024 ( 0.867, 0.961) 72 1.000
0.901 0.040 ( 0.815, 0.965) 156 1.000
0.982 0.015 ( 0.943, 0.999) 65 1.000
$\sigma$ 0.690 0.099 ( 0.519, 0.911) 105 1.000
0.923 0.130 ( 0.696, 1.2) 175 1.000
0.196 0.055 ( 0.103, 0.32) 79 1.000
$\rho$0.544 0.139 (-0.79, -0.253) 116 0.000
0.062 0.132 (-0.203, 0.313) 151 0.672
-0.305 0.130 (-0.549, -0.038) 41 0.012
$\beta$ 0.651 0.074 ( 0.507, 0.797) 45 1.000
0.735 0.075 ( 0.589, 0.881) 52 1.000
0.440 0.057 ( 0.328, 0.552) 12 1.000
$h_{100}$0.680 0.957 (-2.692, 1.076) 56 0.242
-5.493 1.007 (-7.454, -3.467) 72 0.000
2.873 0.358 ( 2.208, 3.634) 17 1.000
Table 5: Posterior mean, standard deviation, 95% credible interval, inefficient factor and the posterior probability that the parameter is positive. TB (top row), DGS (middle row), and AAA (bottom row).
Table 5 shows posterior means, standard deviations, 95% credible intervals, inefficiency factors for parameters (IF), and the posterior probability that the parameter is positive for three datasets. Further, the IF’s for log volatilities, $h_{t}$’s are found to be less than 80, which implies that our mixture sampler is highly efficient as shown in Section 4. The coefficient $\beta$ is estimated to be greater than 0.4, and we find a strong evidence of the positive risk premium since the posterior probability that $\beta$ is positive is almost one, $Pr(\beta>0|y)\approx 1.000$. The autoregressive parameter $\phi$ for the log volatility process is estimated to be more than 0.9, suggesting the high persistence in the volatility as found in the various empirical studies in the previous literature. The correlation parameter $\rho$ is estimated to be negative for TB data and AAA data, implying a strong evidence of the leverage effect since the posterior probability that $\rho$ is negative is almost one, $Pr(\rho<0|y)\approx 1.000$. On the other hand, there is no evidence that $\rho$ is negative for DGS data. Finally, Figures 9, 10 and 11 show 95% credible intervals and posterior medians for $h$ in the case of TB, DGS and AAA data. Noting that
@@
we further plotted moving average $\sum_{s=-10}^{10}z_{t+s}/21$ for reference where $z_{t}=\log(y_{t}^{2})-E(\log\chi_{1}^{2}(\beta^{2}))$ is evaluated at the posterior means of $\beta$. The expected values of $\log\chi_{1}^{2}(\beta^{2})$ are computed numerically as $-0.88,-0.78$, and $-1.08$ for TB, DGS, and AAA using Monte Carlo integration. The traceplot of the estimated log volatilities is similar to that of the moving average series taking account of 95% credible intervals. The large volatilities around years 1980, 2008 and 2010-2023 are well captured by the proposed model as shown in Figures 9, 10 and 11 respectively.
@@
Figure 11: Log volatilities: Moving average of $\log(y_{t}^{2})-E(\log\chi_{1}^{2}(\beta^{2}))$, 95% credible intervals and posterior median. AAA data.
@@ 6.3 Model comparison
In this section we use Bayesian marginal likelihoods to conduct a comprehensive comparisons of various volatility models that includes GARCH models. We calculate the marginal likelihood using the method of Chib (1995). Suppressing the model index, this method is based on an identity introduced in that paper:
@@
where the first term on the right side is the log of the likelihood. the second term is the prior, and the third is the posterior density. We evaluate each of these terms with the posterior mean of $\theta$. For each model, we calculate the first term using the particle filter method given in Section 3.2, setting $I=80,000$. To compute the posterior density ordinate, we apply Chib and Jeliazkov (2001) to the MCMC draws from Algorithm 4 in Appendix C. As Engle et al. (1987) considered three volatility in mean models in the context of the autoregressive conditional heteroscedasticity in mean (ARCH-M) model. The GARCH-M-type models are defined as
$y_{t}$ $=\beta\varphi(h_{t})+u_{t},\quad u_{t}=\sqrt{h_{t}}\nu_{t},$ (18)
$h_{t}$ $=\omega+\gamma h_{t-1}+\alpha u_{t-1}^{2},\quad h_{1}=\frac{\omega}{1-\alpha-\gamma},$ (19)
[…]
  • GARCH-M model: $\varphi(h_{t})=\sqrt{h_{t}}$.
  • GARCH-M $\mathrm{I}\mathrm{I}$ model: $\varphi(h_{t})=h_{t}$.
  • GARCH-M $\mathrm{I}\mathrm{I}\mathrm{I}$ model: $\varphi(h_{t})=\log h_{t}$.
and GARCH model (GARCH, $\beta=0$). The EGARCH-M models are also defined by replacing (19) with
$\log h_{t}=\omega+\gamma\log h_{t-1}+\alpha\nu_{t-1}+\delta(|\nu_{t-1}|-E[|\nu_{t-1}|]),\quad\log h_{1}=\frac{\omega}{1-\gamma},$
where $|\gamma|<1$. As in GARCH-M models, we consider EGARCH-M, EGARCH-M $\mathrm{I}\mathrm{I}$, EGARCH-M $\mathrm{I}\mathrm{I}\mathrm{I}$, EGARCH models for the comparison. The prior distributions are chosen to reflect no prior information regarding of model parameters. Correspondingly, we additionaly consider
  • SVM $\mathrm{I}\mathrm{I}$ model: $y_{t}=\beta\exp(h_{t})+\epsilon_{t}\exp(h_{t}/2)$.
  • SVM $\mathrm{I}\mathrm{I}\mathrm{I}$ model: $y_{t}=\beta h_{t}+\epsilon_{t}\exp(h_{t}/2)$.
@@
Table 6 presents the logarithm of the marginal likelihood for TB, DGS, and AAA data. As for TB and DGS data, the SVML model yields the largest log marginal likelihood among those competing models, which implies our proposed model best describes the risk premium and the time varying volatility among competing models. However, taking account of the standard error, the SVM model performs as well as SVML model for DGS data. Note that our proposed models (SVM and SVML) outperform other SV and SVM models. These results are also consistent with high posterior probabilities of $Pr(\beta>0|y)$ for TB and DGS data, and of $Pr(\rho<0|y)$ for TB data as given in Table 5 of Section 6.2. In most cases, the class of SV models outperforms that of the GARCH and EGARCH models. This comparison shows that the SVM-type models fit better than the GARCH-M and EGARCH-M-type models in the analysis for TB and DGS data and that our proposed SVM model outperform other comprehensive volatility in mean models.
On the other hand, for AAA data, we found the GARCH-M $\mathrm{I}\mathrm{I}\mathrm{I}$ model attains the largest marginal likelihood. This is somewhat surprising, but it may be because the length of the bond holding period is much longer than those of TB and AAA data. Moreover, the volatility clustering in Figure 8 looks different from those in Figures 6 and 7 for TB and DGS data. Also, the level of the volatilities in Figure 11 changes more smootherly than those in Figures 9 and 10. However, we note that the SVML model still outperform other SV models.
Model TB DGS AAA
SVM -185.636(0.156) 66.390(0.109) -1131.540(0.051)
SVML -178.409(0.111) 66.616(0.122) -1126.976(0.061)
SVM $\mathrm{I}\mathrm{I}$211.910(0.072) 52.084(0.046) -1136.338(0.047)
SVML $\mathrm{I}\mathrm{I}$203.926(0.064) 49.823(0.046) -1131.055(0.066)
SVM $\mathrm{I}\mathrm{I}\mathrm{I}$215.386(0.049) 38.438(0.055) -1141.540(0.041)
SVML $\mathrm{I}\mathrm{I}\mathrm{I}$205.118(0.052) 39.304(0.046) -1139.085(0.045)
SV -224.784(0.090) 21.126(0.122) -1155.443(0.049)
SVL -214.247(0.195) 22.549(0.060) -1154.006(0.153)
GARCH-M -215.283(0.003) 10.194(0.005) -1125.723(0.007)
GARCH-M $\mathrm{I}\mathrm{I}$238.244(0.006) 1.940(0.005) -1129.148(0.010)
GARCH-M $\mathrm{I}\mathrm{I}\mathrm{I}$233.935(0.133) 14.178(0.231) -1119.157(0.161)
GARCH -259.238(0.011) -21.886(0.009) -1165.468(0.010)
EGARCH-M -229.199(0.043) 5.297(0.003) -1144.959(0.006)
EGARCH-M $\mathrm{I}\mathrm{I}$244.028(0.009) 4.743(0.007) -1151.550(0.006)
EGARCH-M $\mathrm{I}\mathrm{I}\mathrm{I}$253.988(0.010) -2.3118(0.005) -1143.756(0.005)
EGARCH -254.035(0.005) -14.376(0.004) -1168.813(0.008)
@@
Table 6: Log marginal likelihood estimation and standard error (in parentheses). TB, DGS, and AAA data. SVML, $\mathrm{I}\mathrm{I}$ and $\mathrm{I}\mathrm{I}\mathrm{I}$, and SVL model include the leverage effect $\rho$. The bold font indicates the largest marginal likelihood.
@@ 7 Conclusion
In this paper, we have successfully extended the mixture sampler for the SV model to the SVM model of which the mean equation is described by the standard deviation of the error term as an independent variable. Our main point is the approximation of the distribution of $\log\chi_{1}^{2}(\beta^{2})$ by mixture of normal distributions which is dependent on the parameter $\beta$. This approximation facilitates efficient sampling, leveraging well-established methods for the linear Gaussian state-space model. It is shown in simulation studies that our proposed method is implemented easily and works efficiently. In the empirical studies of the excess holding yield data, we conducted the extensive model comparison including the SVM, GARCH-M and EGARCH-M models, and found that our proposed SVM outperforms other models in terms of the marginal likelihood. It also shows that there exists the positive risk premiums and time-varying volatility in the excess holding yield data.
@@ Acknowledgement
This work is partially supported by JSPS KAKENHI [Grant number: 24H00142]. The computational results are obtained using Rcpp and Ox (see Doornik (2007)).
@@ Appendix A MH step to correct the approximation error
Algorithm 2. Let us denote $\theta=(\alpha,\beta)$ where $\alpha=(\mu,\phi,\sigma^{2})$. The Markov chain Monte Carlo simulation is implemented in four blocks:
@@
$\mu$ 0 0.109 0.326 (-0.616, 0.756) 31 0.103 0.305 (-0.479, 0.802) 84
$\phi$ 0.97 0.971 0.010 ( 0.949, 0.989) 23 0.971 0.011 ( 0.945, 0.990) 110
$\sigma$ 0.3 0.265 0.035 ( 0.202, 0.339) 15 0.263 0.035 ( 0.199, 0.343) 139
$\beta$ 0.7 0.704 0.037 ( 0.631, 0.776) 5 0.731 0.037 ( 0.659, 0.804) 37
$h_{250}$ 2.310 1.881 0.454 ( 1.007, 2.792) 4 1.813 0.457 ( 0.966, 2.722) 74
$h_{750}$ 2.077 1.662 0.404 ( 0.901, 2.493) 5 1.660 0.399 ( 0.924, 2.441) 69

### v2

@@ 6.2 Estimation results for SVM model
Using the same prior distributions as in illustrative examples in Section 4, the proposed SVM models with leverage are fitted to TB, DGS and S&P500 data. We iterated MCMC simulation 50,000 times after discarding initial 10,000 MCMC draws as burn-in period using Algorithm 3 in Appendix B. The acceptance rates of the MH algorithms for $\alpha$ and $(\alpha,h)$ are 72.0% and 20.8% with TB data, 69.5% and 12.8% with DGS data, and 77.3% and 58.9% with S&P500 data, respectively.
@@
$\mu$1.817 0.570 (-2.941, -0.698) 57 0.000
-3.665 0.635 (-4.763, -2.217) 73 0.000
-0.039 0.215 (-0.474, 0.382) 10 0.425
$\phi$ 0.920 0.024 ( 0.865, 0.960) 22 1.000
0.903 0.038 ( 0.813, 0.962) 93 1.000
0.950 0.011 ( 0.927, 0.969) 14 1.000
$\sigma$ 0.685 0.097 ( 0.512, 0.894) 18 1.000
0.911 0.120 ( 0.698, 1.179) 46 1.000
0.327 0.037 ( 0.260, 0.404) 21 1.000
$\rho$0.546 0.139 (-0.784, -0.252) 87 0.000
0.051 0.129 (-0.207, 0.305) 37 0.659
-0.698 0.062 (-0.805, -0.563) 11 0.000
$\beta$ 0.649 0.073 ( 0.507, 0.793) 29 1.000
0.734 0.076 ( 0.587, 0.884) 47 1.000
0.060 0.032 (-0.003, 0.124) 5 0.969
$h_{100}$0.648 0.977 (-2.744, 1.108) 18 0.259
-5.490 1.018 (-7.498, -3.576) 49 0.000
-2.252 0.511 (-3.251, -1.248) 4 0.000
Table 7: Posterior mean, standard deviation, 95% credible interval, inefficient factor and the posterior probability that the parameter is positive. TB (top row), DGS (middle row), and S&P500 (bottom row).
Table 7 shows posterior means, standard deviations, 95% credible intervals, inefficiency factors for parameters (IF), and the posterior probability that the parameter is positive for three datasets. Further, the IF’s for log volatilities, $h_{t}$’s are found to be less than 80, which implies that our mixture sampler is highly efficient as shown in Section 4. The coefficient $\beta$ is estimated to be greater than 0.6 for TB and DGS data, albeit close to zero for S&P500 data. In all cases, we find strong evidence of the positive risk premium since the posterior probability that $\beta$ is positive is almost one for TB and DGS data and 0.97 for S&P500 data. The autoregressive parameter $\phi$ for the log volatility process is estimated to be more than 0.9, suggesting the high persistence in the volatility as found in the various empirical studies in the previous literature. The correlation parameter $\rho$ is estimated to be negative for TB data and S&P500 data, implying a strong evidence of the leverage effect since the posterior probability that $\rho$ is negative is almost one, $Pr(\rho<0|y)\approx 1.000$. On the other hand, there is no evidence that $\rho$ is negative for DGS data. Finally, Figures 9, 10 and 11 show 95% credible intervals and posterior medians for $h$ in the case of TB, DGS and S&P500 data. Noting that
@@
we further plotted moving average $\sum_{s=-10}^{10}z_{t+s}/21$ for reference where $z_{t}=\log(y_{t}^{2})-E(\log\chi_{1}^{2}(\beta^{2}))$ is evaluated at the posterior means of $\beta$. The expected values of $\log\chi_{1}^{2}(\beta^{2})$ are computed numerically as $-0.88,-0.78$, and $-1.27$ for TB, DGS, and S&P500 using Monte Carlo integration. The traceplot of the estimated log volatilities is similar to that of the moving average series taking account of 95% credible intervals. The large volatilities around years 1980, 2008 and 2020/3-2020/4 are well captured by the proposed model as shown in Figures 9, 10 and 11 respectively.
@@
Figure 11: Log volatilities: Moving average of $\log(y_{t}^{2})-E(\log\chi_{1}^{2}(\beta^{2}))$, 95% credible intervals and posterior median. S&P500 data.
@@ 6.3 Model comparison
In this section we use Bayesian marginal likelihoods to conduct a comparison of different stochastic volatility models. We calculate the marginal likelihood using the method of Chib (1995). Suppressing the model index, this method is based on an identity introduced in that paper:
@@
where the first term on the right side is the log of the likelihood. the second term is the prior, and the third is the posterior density. We evaluate each of these terms with the posterior mean of $\theta$. For each model, we calculate the first term using the particle filter method given in Section 3.2, setting $I=80,000$. To compute the posterior density ordinate, we apply Chib and Jeliazkov (2001) to the MCMC draws from Algorithm 4 in Appendix C.
@@
The SVML model had the highest log marginal likelihood for TB and S&P500 data, while the SVM model had the highest for DGS data. This implies our proposed model best describes the risk premium and the time varying volatility among competing models including the standard SV and SVL models. These results are also consistent with high posterior probabilities of $Pr(\beta>0|y)$ for TB, DGS and S&P500 data, and of $Pr(\rho<0|y)$ for TB and S&P500 data as given in Table 7 of Section 6.2.
Model TB DGS S&P500
SVM -186.132(0.019) 68.422(0.016) -1533.027(0.075)
SVML -181.449(0.020) 66.695(0.015) -1508.296(0.059)
SV -231.392(0.020) 16.954(0.020) -1536.776(0.060)
SVL -223.653(0.032) 16.150(0.021) -1509.656(0.039)
@@
Table 8: Log marginal likelihood estimation and standard error (in parentheses). TB, DGS, and S&P500 data. SVML and SVL models include the leverage effect $\rho$. The bold font indicates the largest marginal likelihood.
@@ 7 Conclusion
In this paper, we have successfully extended the mixture sampler for the SV model to the SVM model of which the mean equation is described by the standard deviation of the error term as an independent variable. Our main point is the approximation of the distribution of $\log\chi_{1}^{2}(\beta^{2})$ by mixture of normal distributions which is dependent on the parameter $\beta$. This approximation facilitates efficient sampling, leveraging well-established methods for the linear Gaussian state-space model. It is shown in simulation studies that our proposed method is implemented easily and works fast and efficiently. In the empirical studies of the excess holding yield data and S&P500 data, we conducted the model comparison among the SV and SVM models with and without leverage, and found that our proposed SVM models outperform other models in terms of the marginal likelihood. It shows that there exists the positive risk premiums and time-varying volatilities for all data, while the leverage effects are found to exist for TB and S&P500 data.
@@ Acknowledgement
The authors thank the editor and anonymous referees for their helpful comments. This work is partially supported by JSPS KAKENHI [Grant number: 24H00142]. The computational results are obtained using Rcpp and Ox (see Doornik (2007)).
@@ Appendix A MH step to correct the approximation error
Algorithm 2 (GMS+MH algorithm, GMH). Let us denote $\theta=(\alpha,\beta)$ where $\alpha=(\mu,\phi,\sigma^{2})$. The Markov chain Monte Carlo simulation is implemented in four blocks:
@@
$\mu$ 0 0.114 0.304 (-0.510, 0.703) 5 0.103 0.319 (-0.514, 0.744) 90
$\phi$ 0.97 0.971 0.010 ( 0.948, 0.988) 6 0.971 0.011 ( 0.947, 0.988) 78
$\sigma$ 0.3 0.266 0.036 ( 0.202, 0.342) 9 0.262 0.035 ( 0.196, 0.338) 177
$\beta$ 0.7 0.704 0.037 ( 0.633, 0.776) 3 0.732 0.037 ( 0.661, 0.805) 43
Furthermore, MCMC estimation results using Chan’s method are shown in Tables 12, 13 and for $\beta=0.3,0.5$ and 0.7. Our proposed methods are found to be more efficient with respect to model parameters as in sampling $h_{t}$’s.
CHN $(\beta=0.3)$
Par True Mean Std Dev 95% interval IF
$\mu$ 0 0.087 0.300 (-0.516, 0.680) 27
$\phi$ 0.97 0.971 0.011 ( 0.947, 0.989) 174
$\sigma$ 0.3 0.262 0.037 ( 0.196, 0.341) 249
$\beta$ 0.3 0.328 0.033 ( 0.263, 0.393) 12
CHN $(\beta=0.5)$
Par True Mean Std Dev 95% interval IF
$\mu$ 0 0.088 0.319 (-0.573, 0.726) 18
$\phi$ 0.97 0.972 0.011 ( 0.949, 0.991) 141
$\sigma$ 0.3 0.258 0.036 ( 0.192, 0.339) 266
$\beta$ 0.5 0.530 0.034 ( 0.464, 0.600) 23
CHN $(\beta=0.7)$
Par True Mean Std Dev 95% interval IF
$\mu$ 0 0.095 0.312 (-0.527, 0.722) 25
$\phi$ 0.97 0.971 0.010 ( 0.950, 0.988) 173
$\sigma$ 0.3 0.262 0.034 ( 0.197, 0.331) 277
$\beta$ 0.7 0.730 0.037 ( 0.658, 0.803) 126